Use the quadratic formula to solve each of the following quadratic equations.
2x^2-5x=3

Answer by checkley75(3666)

2x^2-5x=3
2x^2-5x-3=0
(2x+1)(x-3)=0
2x+1=0
2x=-1
x=-1/2 answer
x-3=0
x=3 answer.